Corporate Finance Core Principles and Applications 5th Edition by Stephen Ross (1).pdf - Adobe Acrobat Reader DC File Edit View Window Help Home Tools Corporate Finance X Sign In 0 Θ T 也Share 229 (262 of 721) 104%. At least one signature has problems. Signature Panel Export PDF Seth Bullock, the owner of Bullock Gold Mining, is eveluating a new gold mine in South Dakota. Dan Dority, the companys geologist, has just firished his analysis of the mine site. He has estimated that the mine would be productive for eightyears, after which the gold would be completely mined. Dan has taken an estimate of the gold deposits to Alma Garrett, the companys financial officer. Alma has been asked by Seth to perform an analysis of the new mine and present her recommendation on whether the company should open the new mine. Adobe Export PDF Convert PDF Files to Word or Excel Online Select PDF File Alma has used the estimates provided by Dan to determine the revenues that could be expected from the mine. She has aso projected the expense of opening the mine and the annual operating expenses. If the company opens the mine, it will cost $850 million today, and it will have a cash outlow of $120 million nine years from today in costs associated with closing the mine and reclaiming the area surounding it. The expected cash flows each year from the mine are shown in the table that follows. Bullock has a 12 percent required return on all of its gold mines. Corporate Ss (1)pdf X Convert to LO Microsoft Word (docx) YEAR CASH FLOW Document Language: English (U.S) Change -$850,000,000 165,000,000 190,000,000 225,000,000 245,000,000 235,000,000 195,000,000 175,000,000 155,000,000 -120,000,000 Convert Create PDF v Store and share files in the Document Cloud 1. Construct a spreadsheet to calculate the payback period, internal rate of returm, modified internal rate of return, and net present value of the p roposed mine Learn More 2. Based on your analysis, should the company open the mine? 4:12 AM 1/27/2019 O Type here to search aR ^ ョ塩/
Answer #1
NPV: 0 1 2 3 4 5 6 7 8 9
Cash flow -850000 165000 190000 225000 245000 235000 195000 175000 155000 -120000
PVIF at 12% 1 0.89286 0.79719 0.71178 0.63552 0.56743 0.50663 0.45235 0.40388 0.36061
PV at 12% -850000 147321 151467 160151 155702 133345 98793 79161 62602 -43273
NPV 95269
IRR:
IRR is that discount rate for which NPV = 0. It has to be found out by trial and error. Different interest rates are to be
used till 0 NPV is reached.
Discounting with 15% NPV
PVIF at 15% 1 0.86957 0.75614 0.65752 0.57175 0.49718 0.43233 0.37594 0.32690 0.28426
PV at 15% -850000 143478 143667 147941 140080 116837 84304 65789 50670 -34111 8654
Discounting with 16%
PVIF at 16% 1 0.86207 0.74316 0.64066 0.55229 0.47611 0.41044 0.35383 0.30503 0.26295
PV at 16% -850000 142241 141201 144148 135311 111887 80036 61920 47279 -31554 -17531
IRR lies between 15% and 16% as 0 NPV will fall between these discount rates.
By simple interpolation, IRR = 15+8654/(8654+17531) = 15.33
MIRR:
MIRR assumes reinvestment of intermediate cash flows at the WACC of 12%. The Cumulative FV of all cash flows at t9 is to be found out.
FVIF at 12% 2.47596 2.21068 1.97382 1.76234 1.57352 1.40493 1.25440 1.12000 1.00000
FV at 12% of CFs t1 to t9 408534 420029 444110 431774 369777 273961 219520 173600 -120000 2621305
MIRR = (2621305/850000)^(1/9)-1 = 13.33%
PAYBACK PERIOD:
Cumulative cash flows (including the
mining restoration costs are t9) -970000 -805000 -615000 -390000 -145000 90000 285000 460000 615000
Payback period = 4+145000/235000 = 4.62 Years